A rectangle has a length that is 2less than 3times the width if the area of the rectangle is 16cm²,find the the dimentions

Answer by fractalier(6550)

Call the width, W.
Then the length, L, is found by
L = 3W - 2
The area A is found by
A = LW = 16
Now substitute the first into the second and get
(3W-2)W=16
3W^2-2W-16=0
Solve by factoring...
(3W-8)(W+2)=0
so that 3W-8 =0
and W = 8/3 cm
which leads to L = 6 cm